# Blockchain-Enabled Privacy-Preserving Second-Order Federated Edge Learning in Personalized Healthcare

### v1

@@ Abstract
Federated learning (FL) has attracted increasing attention to mitigate security and privacy challenges in traditional cloud-centric machine learning models specifically in healthcare ecosystems. FL methodologies enable the training of global models through localized policies, allowing independent operations at the edge clients’ level. Conventional first-order FL approaches face several challenges in personalized model training due to heterogeneous non-independent and identically distributed (non-iid) data of each edge client. Recently, second-order FL approaches maintain the stability and consistency of non-iid datasets while improving personalized model training. This study proposes and develops a verifiable and auditable optimized second-order FL framework BFEL (blockchain enhanced federated edge learning) based on optimized FedCurv for personalized healthcare systems. FedCurv incorporates information about the importance of each parameter to each client’s task (through Fisher Information Matrix) which helps to preserve client-specific knowledge and reduce model drift during aggregation. Moreover, it minimizes communication rounds required to achieve a target precision convergence for each edge client while effectively managing personalized training on non-iid and heterogeneous data. The incorporation of Ethereum-based model aggregation ensures trust, verifiability, and auditability while public key encryption enhances privacy and security. Experimental results of federated CNNs and MLPs utilizing Mnist, Cifar-10, and PathMnist demonstrate the high efficiency and scalability of the proposed framework.
Index Terms:
Federated Learning, FedCurv, Data Privacy, Blockchain, Personalized Healthcare, Federated Edge Learning, Second-Order Federated Learning
@@ I Introduction
Traditional machine learning (ML) methodologies necessitate training on data consolidated within a single data repository, which may be either centralized or distributed [1]. It requires raw data from multiple participants to be transmitted to a centralized server (where the data are aggregated). However, aggregating data from multiple stakeholders in healthcare systems poses significant challenges, particularly concerning security, compromising data owners’ privacy, and possibly exposing sensitive health data [2].
Within the distributed ML paradigm, two primary frameworks exist: data centre-based distributed ML and cross-device FL. The former utilizes optimized computing nodes, data shuffling, and high-bandwidth communication networks, whereas the latter operates on a large number of resource-constrained devices with limited computational, storage, and communication capabilities [3].
@@
Federated edge learning (FEL) has emerged as a distributed machine learning paradigm that mitigates privacy concerns by facilitating collaborative model training while ensuring that data remains decentralized. FEL bridges these two paradigms, possessing computational and storage capacities comparable to data centre-based ML while sharing communication constraints with cross-device FL due to physical distance between edge nodes, multi-hop transmissions, and diverse communication mediums. It offers a decentralized alternative, enabling multiple devices to collaboratively train a model without sharing raw data, where each participant trains a model locally, and shares only the model parameters. Some of the widely used gradient-based first-order federated learning approaches are FedAvg, FedSgd, FedAdam, and FedYogi. While these methodologies preserve the confidentiality of the sensitive data, the resulting shared model parameters remain vulnerable to confidentiality breaches during aggregation and dissemination. Despite its several advantages, FL faces two fundamental challenges: (i) managing heterogeneous systems within the federated network and (ii) handling real-world data that are often non-independently and identically distributed (non-iid) among clients [4]. Specifically, the delivery of personalized FL services stands crucial in healthcare advancement because patients manifest individual health profiles with specific requirements. The implementation of first-order FL algorithms encounters substantial hurdles when it comes to personalizing deep learning and machine learning models [5]. Such training datasets have limited representation of particular classes or behaviors because specific data points are spread sparsely throughout the dataset.
@@
However, second-order FL methods provide better suitability for varied heterogeneous healthcare data present in different edge client devices. It allows better customization of local models thus enhancing their value for personalized healthcare systems. Such adaptations provide potential benefits of privacy and security in heterogeneous environments while preserving personalization (i.e. higher model adaption variability) such as Natural Gradient Descent [6] and Quasi-Newton method [7]. Nevertheless, the distributed implementation of second-order methodologies in a traditional manner remains challenging due to their reliance on inverse matrix-vector product computations, which introduce complexity in determining the descent direction. Addressing these challenges, it is crucial to develop efficient and scalable second-order optimization techniques tailored to heterogeneous FL environments. Second-order optimization methods offer a key advantage by incorporating curvature information of the loss function, thus improving personalized training processes.
@@
This study aims to optimize a distributed optimization algorithm FedCurv (Federated Curvature) that minimizes communication rounds required to achieve a target precision for convergence while effectively managing personalized training on non-iid and heterogeneous data across edge client devices. The main contributions of this article:
@@
    We proposed optimized implementation of FedCurv for heterogeneous medical data to improve personalized health monitoring and prediction.
@@
    Blockchain-based secure aggregation mechanisms to ensure trust, tamper-proof model updates, decentralized auditability, and accountability.
@@
The remainder of this article is organized as follows. Section II presents the background and motivation, Section III details the proposed BFEL framework, and Section IV describes the experimental setup. Section V discusses the results, and Section VI concludes the paper with future research directions.
@@ II Related Work
In recent years, advancements in AI, particularly in the domain of DL within ML, have significantly contributed to the enhancement of smart and personalized healthcare ecosystems [10]. This progress has been largely driven by the expansion of training datasets. However, as the volume of training data increases, so too does the associated risk of privacy breaches [11]. Studies indicate that privacy attacks targeting DL models can result in the inadvertent leakage of sensitive training data. Such privacy concerns present a substantial barrier to the continued advancement and deployment of DL technologies [12]. Centralized aggregation of private data for ML models poses significant challenges, particularly in terms of security, privacy, and confidentiality [13]. Consequently, retaining control over data without external dissemination to prioritize the security and privacy of data owners is an essential requirement for healthcare applications. This requirement has led to the development of FL methodologies. Auditability and verifiability are essential components for establishing trustworthiness in FL. Several studies proposed blockchain-empowered FL approaches to enhance transparency, accountability, verifiability and the independent validation of FL processes [9]. Authors in [14] proposed a blockchain based trusted execution platform to secure each client’s local model training data and provide multi-signature-based global model verification to enhance model auditability. In [15] authors develop smart contract empowered local data training policy control and verification of the integrity of trained models.
@@
In the literature, first-order federated edge learning techniques, which rely solely on gradient information, are frequently utilized due to their robustness in distributed settings and minimal local computational requirements [16]. Conversely, second-order methods, utilize both gradient and curvature information, thereby facilitating improved descent direction selection and significantly accelerating convergence. This acceleration reduces the number of communication rounds required to achieve convergence, making second-order methods particularly advantageous in heterogeneous FL environments.
@@
Continual learning and FL approaches employ diverse strategies to address challenges like catastrophic forgetting, task interference, and communication efficiency [17]. Elastic Weight Consolidation (EWC) [18] mitigates catastrophic forgetting by restricting parameter updates critical to previous tasks using FIM, ensuring solutions compatible with both old and new tasks. In contrast, Incremental Moment Matching (IMM) [19] models the posterior distribution of parameters for multiple tasks as a mixture of gaussians to harmonize task-specific knowledge. Stable SGD [20] enhances performance by dynamically adjusting hyperparameters and incrementally reducing the learning rate upon encountering new tasks. For FL, FedCurv [21] adapts a modified EWC framework to minimize disparities between client models during collaborative training. Recent advancements [22] further refine aggregation mechanisms using Bayesian non-parametric methods to improve model alignment across heterogeneous data sources. Communication overhead, a persistent hurdle in federated systems, is addressed [23], which employs layer-wise aggregation where shallow layers are updated frequently, while deeper layers are consolidated only in the final stages of training loops, significantly reducing bandwidth demands. These methods collectively advance the robustness and scalability of learning systems in dynamic, distributed environments.
@@ III Privacy preserving BFEL Service Framework
This section provides an architectural flow of a proposed privacy-preserving blockchain enhanced federated edge learning (BFEL) framework, which maintains auditability, verifiability, availability and ensures privacy protection in edge-FL environments. Second-order federated edge learning based optimized Federated Curvature (FedCurv) algorithm is proposed for heterogeneous smart personalized healthcare ecosystems. FedCurv algorithm is built on Elastic Weight Consolidation (EWC) to prevent catastrophic forgetting across edge clients and incorporates second-order information from the Fisher Information Matrix (FIM) to preserve critical model parameters during the training process. Figure 2 depicts the system level flow of the proposed BFEL system.
@@ III-A Edge Client Layer
The proposed architecture shown in Figure 3 contains the first tier, namely, the edge clients layer. This layer operates through edge devices, including wearable technologies and Internet of Things (IoT) sensors, which acquire real-time health metrics. Data sources encompass three primary categories: patients utilizing personal health monitoring devices, hospital electronic health record (EHR) systems, and diagnostic datasets from clinical or laboratory settings. In a network of $n$ clients denoted by $k=\left\{k_{1},k_{2},\ldots,k_{n}\right\}$ each hosting $m$ IoT devices for data acquisition, $D=\left\{d_{1},d_{2},\ldots,d_{m}\right\}$, we consider. The service run by each IoT device $d_{j}\in D$, referred to as the collection service ($S_{\text{col }}$) is responsible for gathering data (e.g. EEG signals).
@@
Following this, local training is performed with curvature regularization. The goal is to minimize a regularized loss function $\mathcal{L}_{k}(\theta)$, which combines the standard local loss with a curvature penalty term. The penalty term involves the squared difference between local and global parameters weighted by the diagonal FIM. The gradient of the regularized loss is then computed as the sum of the gradient of the local loss and a regularization term scaled by a factor $\lambda$. The model parameters are updated using stochastic gradient descent (SGD) for $E$ local epochs based on the gradient of this regularized loss. $F_{k}[i]$ corresponds to diagonal entry of $F_{k}$ for parameter $i$.
@@
above equation shows the gradient of the log-likelihood with respect to parameters.
Compute local training with curvature regularization to minimize the regularized loss:
@@
Subsequently, update $\theta$ via SGD for $E$ epochs:
@@
In Post-training, clients transmit $\theta_{\text{local}}$ curvature and $g_{k}$ gradients to the aggregation server through blockchain. This approach inherently safeguards data privacy by ensuring sensitive information remains within its original jurisdiction, serving as the primary defence against privacy breaches in sensitive healthcare applications. These updates are systematically broadcast to a private blockchain network ensuring auditability and transparency.
@@ III-B Blockchain Layer
The blockchain layer serves as a foundational component of the proposed system, providing an immutable and transparent ledger to enhance security, accountability, and compliance in healthcare ecosystems. By design, the blockchain records all transactions including consent management, data access requests, and model updates in a tamper-evident manner, ensuring traceability across the ecosystem.
For instance, patient consent parameters, such as permissible data usage and authorized entities, are codified within smart contracts. These self executing agreements autonomously verify compliance with predefined policies, thereby eliminating manual oversight and reducing the risk of unauthorized access. Each step of the aggregation workflow such as server update requests, participant submissions, and the generation of the global model is cryptographically hashed and immutably logged on the blockchain, enabling independent verification of procedural integrity.
Bloackchain broadcasts stored the finalized global model parameters in encrypted form using public-key infrastructure $pki$, ensuring only authorized entities with corresponding decryption keys can access sensitive insights. Subsequent training rounds are initiated by broadcasting these encrypted parameters via the blockchain to aggregation servers, thereby maintaining synchronization while preserving security. Access control is further reinforced through blockchain mediated authentication; for example, researchers requesting model access trigger smart contracts that validate their permissions against patient defined policies prior to granting approval. All such interactions, including consent modifications and data retrieval attempts, are permanently recorded on the ledger, creating auditable trails for regulatory review. This architecture not only ensures adherence to standards such as HIPAA and GDPR but also empowers patients with visibility into data usage through transparent audit logs. By integrating decentralized consensus mechanisms, cryptographic encryption, and automated policy enforcement, the system establishes a robust framework for privacy preserving collaboration, balancing operational transparency with uncompromising data security.
@@ III-C Cloud Server Layer
Cloud servers are deployed virtually over the network and working as aggregation servers, collecting locally computed $\theta_{\text{local}}$ and gradients $g_{k}$ from each edge client $k$. Compute aggregated curvature $F_{\text{global}}$ to get important scores (FIMs) from all clients and average gradients $g_{\text{global}}$.
@@
Finally, calculated $\theta_{\text{global}}^{\text{new}}$ global updates are broadcast to the clients through blockchain broadcasts using smart contracts.

### v2

# Blockchain-Enabled Privacy-Preserving Second-Order Federated Edge Learning in Personalised Healthcare

@@ Abstract
Federated learning (FL) is increasingly recognised for addressing security and privacy concerns in traditional cloud-centric machine learning (ML), particularly within personalised health monitoring such as wearable devices. By enabling global model training through localised policies, FL allows resource-constrained wearables to operate independently. However, conventional first-order FL approaches face several challenges in personalised model training due to the heterogeneous non-independent and identically distributed (non-iid) data by each individual’s unique physiology and usage patterns. Recently, second-order FL approaches maintain the stability and consistency of non-iid datasets while improving personalised model training. This study proposes and develops a verifiable and auditable optimised second-order FL framework BFEL (blockchain enhanced federated edge learning) based on optimised FedCurv for personalised healthcare systems. FedCurv incorporates information about the importance of each parameter to each client’s task (through fisher information matrix) which helps to preserve client-specific knowledge and reduce model drift during aggregation. Moreover, it minimizes communication rounds required to achieve a target precision convergence for each client device while effectively managing personalised training on non-iid and heterogeneous data. The incorporation of ethereum-based model aggregation ensures trust, verifiability, and auditability while public key encryption enhances privacy and security. Experimental results of federated CNNs and MLPs utilizing mnist, cifar-10, and PathMnist demonstrate framework’s high efficiency, scalability, suitability for edge deployment on wearables, and significant reduction in communication cost.
@@ I Introduction
Traditional ML methodologies necessitate training on data consolidated within a single data repository, which may be either centralised or distributed [1]. This paradigm requires raw data from multiple participants to be transmitted to a centralised aggregator server. However, aggregating data from multiple stakeholders in healthcare systems poses significant challenges, particularly concerning security, compromising data owners’ privacy, and possibly exposing sensitive health data and high latency rate [2, 3]. Within the distributed ML paradigm, two primary frameworks exist: data centre-based distributed ML and cross-device FL. The former utilizes optimised computing nodes, data shuffling, and high-bandwidth communication networks, whereas the latter operates on a large number of resource-constrained devices with limited computational, storage, and communication capabilities [4].
@@
Federated edge learning (FEL) has emerged as a distributed ML paradigm that mitigates privacy concerns by facilitating collaborative model training while ensuring that data from wearables remains decentralised on the devices. FEL bridges these two paradigms, possessing computational and storage capacities comparable to data centre-based ML while sharing communication constraints with cross-device FL due to physical distance between devices, multi-hop transmissions, and diverse communication mediums. It offers a decentralised alternative, enabling multiple wearables to collaboratively train a model without sharing raw data, where each participant trains a model locally, and shares only the model parameters. Some of the widely used first-order FL approaches are FedAvg, FedSgd, FedAdam, and FedYogi. While these methodologies preserve the confidentiality of the sensitive data, the resulting shared model parameters remain vulnerable to confidentiality breaches during aggregation and dissemination. Despite its several advantages, FL faces two fundamental challenges: (i)managing hardware and data heterogeneity across a diverse fleet of consumer devices and (ii) handling real-world data that are often non-independently and identically distributed (non-iid) as physiological patterns are unique to each individual [5, 6]. Specifically, the delivery of personalised FL services stands crucial in wearable technologies advancement because patients manifest individual health profiles with specific requirements. Gradient based first-order approaches are widely used in FL to preserve the confidentiality of sensitive local data. However, this protection does not extend to the model parameters themselves, which remain vulnerable to confidentiality breaches during aggregation and dissemination. General training datasets have limited representation of particular classes or behaviors because specific data points are spread sparsely throughout the dataset.  [7].
@@
However, second-order FL methods provide better suitability for varied, heterogeneous wearable data present on different consumer clients. It allows better customisation of local models thus enhancing their value for personalised healthcare systems. Such adaptations provide potential benefits of privacy and security for consumer electronics while preserving personalization (i.e. higher model adaption variability) such as natural gradient descent [8] and quasi-newton method [9].
Nevertheless, the distributed implementation of second-order methodologies in a traditional manner remains challenging due to their reliance on inverse matrix-vector product computations, which introduce complexity in determining the descent direction. Addressing these challenges, it is crucial to develop efficient and scalable second-order optimization techniques tailored to heterogeneous FL environments. Second-order optimization methods offer a key advantage by incorporating curvature information of the loss function, thus improving personalised training processes.
@@
This study aims to optimize a distributed optimization algorithm FedCurv that minimizes communication rounds from resource-limited wearables required to achieve a target precision for convergence while effectively managing personalised training on non-iid and heterogeneous data across consumer edge devices. The main contributions of this article:
@@
    We proposed optimised implementation of FedCurv for heterogeneous medical data to improve real-time personalised health monitoring and prediction for wearable devices.
@@
    Blockchain-based secure aggregation mechanisms to ensure trust, tamper-proof model updates, decentralised auditability, and accountability.
@@
The remainder of this article is organised as follows. Section II presents the background and motivation, Section III details the proposed BFEL framework, and Section IV describes the experimental setup. Section V discusses the results, and Section VI concludes the paper with future research directions.
@@ II Related Work
Recent advancements in AI, particularly in the domain of DL within ML, have significantly contributed to the enhancement of smart and personalised healthcare ecosystems [12]. While the expansion of training datasets has been a key driver of progress, it simultaneously heightens the risk of privacy violations [13]. Research indicates that such violations occur through privacy attacks which can extract sensitive information from the training data. This progress is largely driven by the expansion of training datasets. However, increased data volume also heightens the risk of privacy violations [13]. Studies indicate that privacy attacks targeting DL models can lead to the inadvertent leakage of sensitive training datasets. Such privacy concerns present a substantial barrier to the continued advancement and deployment of DL technologies [14]. Centralised aggregation of private data for ML models poses significant challenges, especially in terms of security, privacy, and confidentiality [15]. Consequently, retaining control over data without external dissemination to prioritize the security and privacy of data owners is an essential requirement for healthcare applications, motivating the creation of FL methods. Auditability and verifiability are essential components for establishing trustworthiness in FL. Several studies proposed blockchain-empowered FL approaches to enhance transparency, accountability, verifiability and the independent validation of FL processes [11].
Achieving trustworthy FL systems requires mechanisms for auditability and verifiability. Numerous studies have suggested combining FL with blockchain technology to improve transparency, accountability, verifiability, and independent auditing of FL operations [11]. For example, some authors designed a blockchain-based trusted execution environment to protect local training data and introduced multi-signature verification for global models to strengthen auditability [16]. Others have created traceable, transparent, and auditable supply chain solutions by optimizing data batching with Hyperledger Sawtooth [17]. Additional work includes smart contract-based local training policy enforcement and integrity verification for trained models [18], as well as peer-to-peer data sharing platforms that use blockchain to return data ownership rights to the original producers [19].
@@
In FL research, first-order federated edge learning techniques, which rely solely on gradient information, are frequently utilised due to their robustness in distributed settings and minimal local computational requirements [20]. In comparison, second-order methods incorporate both gradient and curvature information, thereby facilitating improved descent direction selection and significantly accelerating convergence. This acceleration reduces the number of communication rounds required to achieve convergence, making second-order methods particularly advantageous in heterogeneous FL environments.
@@
Both continual learning and FL approaches employ diverse strategies to address challenges like catastrophic forgetting, task interference, and communication efficiency [21]. Elastic weight consolidation (EWC) [22] mitigates catastrophic forgetting by restricting parameter updates critical to previous tasks using FIM, ensuring solutions compatible with both old and new tasks. In contrast, Incremental moment matching (IMM) [23] models the posterior distribution of parameters for multiple tasks as a mixture of gaussians to harmonize task-specific knowledge. Stable SGD [24] enhances performance by dynamically adjusting hyperparameters and incrementally reducing the learning rate upon encountering new tasks. For FL, FedCurv [25] adapts a modified EWC framework to minimize disparities between client models during collaborative training as depicted in figure 1. Recent advancements [26] further refine aggregation mechanisms using Bayesian non-parametric methods to improve model alignment across heterogeneous data sources. Communication overhead, a persistent hurdle in federated systems, is addressed [27], which employs layer-wise aggregation where shallow layers are updated frequently, while deeper layers are consolidated only in the final stages of training loops, significantly reducing bandwidth demands. Authors in [28] presents FedSecurity, a tool that simplifies testing security attacks and defenses in FL. It saves time by handling the underlying setup, allowing researchers to easily experiment with different models, datasets, and protection methods. In a recent study [29], large language models are trained to work with private data at edge level, allowing owners to collaboratively train the model while utilising parameters from base model. Together, these methods collectively advance the robustness and scalability of learning systems in dynamic, distributed environments.
@@ III Privacy preserving BFEL Service Framework
This section outlines the architectural flow of a proposed privacy-preserving blockchain enhanced federated edge learning (BFEL) framework, which maintains auditability, verifiability, availability and ensures privacy protection in edge-FL environments. To address the challenges of heterogeneous data in smart personalised healthcare devices, we propose an optimised second-order federated learning algorithm named federated curvature (FedCurv). FedCurv algorithm is built on elastic weight consolidation (EWC) to prevent catastrophic forgetting across edge devices and incorporates second-order information from the fisher information matrix (FIM) to preserve critical model parameters during the training process. The system-level flow of the proposed BFEL framework is illustrated in figure 2.
@@ III-A Edge Client Layer
The proposed architecture, shown in figure 3, comprises a multi-tiered system beginning with the edge clients layer. This foundational tier consists of edge devices, such as wearable technologies. Data sources encompass three primary categories: patients utilizing personal health monitoring devices, hospital electronic health record (EHR) systems, and diagnostic datasets from clinical or laboratory settings. In a network of $n$ clients denoted by $k=\left\{k_{1},k_{2},\ldots,k_{n}\right\}$ each hosting $m$ IoT devices for data acquisition, $D=\left\{d_{1},d_{2},\ldots,d_{m}\right\}$, we consider. The service run by each IoT device $d_{j}\in D$, referred to as the collection service ($S_{\text{col }}$) is responsible for gathering data (e.g. EEG signals).
@@
Subsequently, local training is performed using curvature regularization. The objective is to minimize a regularised loss function $\mathcal{L}_{k}(\theta)$, which combines the standard local loss with a curvature penalty term. The penalty term involves the squared difference between local and global parameters weighted by the diagonal FIM. The gradient of the regularised loss is then computed as the sum of the gradient of the local loss and a regularization term scaled by a factor $\lambda$. The model parameters are updated using stochastic gradient descent (SGD) for $E$ local epochs based on the gradient of this regularised loss. $F_{k}[i]$ corresponds to diagonal entry of $F_{k}$ for parameter $i$.
@@
Above equation describes the gradient of the log-likelihood with respect to parameters. Compute local training with curvature regularization to minimize the regularised loss:
@@
Following this, update $\theta$ via SGD for $E$ epochs:
@@
In post-training, clients propagate their local model parameters $\theta_{\text{local}}$, curvature data and $g_{k}$ gradients to the aggregation server through blockchain. This approach inherently safeguards data privacy by ensuring sensitive information remains within its original jurisdiction, serving as the primary defence against privacy breaches in sensitive healthcare applications. These updates are systematically broadcast across a private blockchain network, ensuring auditability and transparency.
@@ III-B Blockchain Layer
The blockchain layer serves as a foundational component of the proposed system, providing an immutable and transparent ledger to enhance security, accountability, and compliance in healthcare ecosystems. By design, the blockchain records all transactions including consent management, data access requests, and model updates in a tamper-evident manner, thereby ensuring full traceability across the network.
Specifically, smart contracts are utilised as patient consent parameters, such as permissible data usage and authorized entities. These self-executing agreements autonomously enforce predefined policies, which eliminates manual oversight and reduces the risk of unauthorized access. Each step of the aggregation workflow such as server update requests, participant submissions, and the generation of the global model is cryptographically hashed and immutably logged on the blockchain, enabling independent verification of procedural integrity.
The bloackchain broadcasts, stored the finalised global model parameters in an encrypted form using public-key infrastructure pkipki, ensuring that only authorised entities with corresponding decryption keys can access sensitive insights. Subsequent training rounds are initiated by broadcasting these encrypted parameters via the blockchain to aggregation servers, thereby maintaining synchronization while preserving security. To further reinforce access control, blockchain mediated authentication ensures that researchers’ access requests trigger smart contracts. These contracts validate permissions against patient-defined policies before any approval is granted. All such interactions, including consent modifications and data retrieval attempts, are permanently recorded on the ledger, creating auditable trails for regulatory review. This architecture not only ensures adherence to standards such as health insurance portability and accountability act (HIPAA) [30] and general data protection regulation (GDPR) [31] but also empowers patients with visibility into data usage through transparent audit logs. By integrating decentralised consensus mechanisms, cryptographic encryption, and automated policy enforcement, the system establishes a robust framework for privacy preserving collaboration, balancing operational transparency with uncompromising data security.
@@ III-C Cloud Server Layer
Cloud servers are deployed virtually over the network and working as aggregation servers, collecting locally computed $\theta_{\text{local}}$ and gradients $g_{k}$ from each client device $k$. Compute aggregated curvature $F_{\text{global}}$ to get important scores (FIMs) from all clients and average gradients $g_{\text{global}}$.
@@
Finally, calculated $\theta_{\text{global}}^{\text{new}}$ global updates are broadcast to the clients through the blockchain broadcasts using smart contracts.